4 Caspian Sea Drinks needs to raise $74.00 million by issuing additional shares of stock. If...

4
Caspian Sea Drinks needs to raise $74.00 million by issuing additional shares of stock. If the market estimates CSD will pay a dividend of $2.03 next year, which will grow at 3.01% forever and the cost of equity to be 11.93%, then how many shares of stock must CSD sell?


Submit
Answer format: Number: Round to: 0 decimal places.


unanswered
not_submitted
#5
Suppose the risk-free rate is 2.36% and an analyst assumes a market risk premium of 5.49%. Firm A just paid a dividend of $1.21 per share. The analyst estimates the β of Firm A to be 1.44 and estimates the dividend growth rate to be 4.09% forever. Firm A has 271.00 million shares outstanding. Firm B just paid a dividend of $1.99 per share. The analyst estimates the β of Firm B to be 0.78 and believes that dividends will grow at 2.30% forever. Firm B has 185.00 million shares outstanding. What is the value of Firm A?


Submit
Answer format: Currency: Round to: 2 decimal places.


unanswered
not_submitted
#6
Suppose the risk-free rate is 1.95% and an analyst assumes a market risk premium of 7.98%. Firm A just paid a dividend of $1.21 per share. The analyst estimates the β of Firm A to be 1.23 and estimates the dividend growth rate to be 4.18% forever. Firm A has 268.00 million shares outstanding. Firm B just paid a dividend of $1.77 per share. The analyst estimates the β of Firm B to be 0.76 and believes that dividends will grow at 2.35% forever. Firm B has 196.00 million shares outstanding. What is the value of Firm B?


Submit
Answer format: Currency: Round to: 2 decimal places.
Answer #1

4). Share Price = D1 / [r - g] = $2.03 / [0.1193 - 0.0301] = $2.03 / 0.0892 = $22.76

Number of Shares = Amount Raised / Share Price = $74 million / $22.76 = 3,251,626 shares

5). According to the CAPM,

Required return = Risk-free Rate + [Beta*Market Risk Premium]

= 2.36% + [1.44*5.49%] = 2.36% + 7.9056% = 10.2656%

Share Price(Firm A) = [D0 * (1 + g)] / [r - g]

= [$1.21 * (1 + 0.0409)] / [0.102656 - 0.0409] = $1.259489 / 0.061756 = $20.39

Value of Firm A = Share Price(Firm A) * Shares Outstanding

= $20.39 * 271 million = $5,526.94 million

6). According to the CAPM,

Required return = Risk-free Rate + [Beta*Market Risk Premium]

= 1.95% + [0.76*7.98%] = 1.95% + 6.0648% = 8.0148%

Share Price(Firm B) = [D0 * (1 + g)] / [r - g]

= [$1.77 * (1 + 0.0235)] / [0.080148 - 0.0235] = $1.811595 / 0.056648 = $31.98

Value of Firm B = Share Price(Firm B) * Shares Outstanding

= $31.98 * 196 million = $6,268.05 million